OpenStudy (anonymous):

. The odds against Lounging Larry winning the horse race are 5:12. Determine the probability that Lounging Larry wins the horse race

OpenStudy (ash2326):

odds against Lounging Larry winning the horse race are 5/12 odds against=Probability of losing/ probability of winning Let P= probability of wining we know that Probability of losing= 1-P do you get this ?

OpenStudy (anonymous):

no ma i really tried youtube and all i dont get it

OpenStudy (ash2326):

Do you agree with this \[P (win)+P(lose)=1\]

OpenStudy (anonymous):

yes thank you so much

OpenStudy (ash2326):

did you find the probability?